OpenStudy (anonymous):

What is the equation of the line in slope-intercept form? the line perpendicular to y = 1/3x+5 through (2, 1) A. y = 1/3x + 7 B. y = -3x + 7 C. y = 3x + 7 D. y = -1/3x + 7

OpenStudy (anonymous):

perpendicular is reciprocal and multiply by -1

OpenStudy (anonymous):

-3 is a line perpendicular to it

OpenStudy (anonymous):

y=mx+b 1=-3*2+b 1=-6+b sfb 7=b so y=-3x+7 B